OpenStudy (anonymous):

convert to decimal notation 10^-3

OpenStudy (anonymous):

It is also equal to \[\frac {1}{1000}\] = ? Can you try now :D

OpenStudy (anonymous):

0.001

OpenStudy (anonymous):

Very good :D

OpenStudy (anonymous):

thnak you

OpenStudy (anonymous):

Welcome. Just one tip 10^2 = 10 * 10 10^3 = 10*10*10 \[10^{-2}=\frac{1}{10}*\frac{1}{10}\] \[10^{-3}=\frac{1}{10}*\frac{1}{10}*\frac{1}{10}\]